x-6= the square root of 3x
Answer by edjones(7311)

x-6=sqrt(3x)
(x-6)^2=3x
x^2-12x+36=3x
x^2-15x+36=0
(x-12)(x-3)=0
x=12, x=3
.
Always check:
12-6=sqrt(3*12)
6=sqrt(36) true
.
3-6=sqrt(3*3)
-3=sqrt(9) false
.
x=12 is the only answer.
.
Ed